OpenStudy (anonymous):

Solve the equation ln(x + 1) - ln(9) = ln(4) + ln(x - 3) for x. A) 2.86 B) 3.11 C) 3.36 D) 3.61 E) 3.86

OpenStudy (anonymous):

use properties of ln: ln((x+1)/9)=ln(4*(x-3)) so, (x+1)/9 = 4*(x-3) simplify & solve for x

OpenStudy (anonymous):

x = 13/4

OpenStudy (anonymous):

= 3.25

OpenStudy (anonymous):

which is none of the answers D:

OpenStudy (anonymous):

I got my equation wrong and didn't look at the options first

OpenStudy (anonymous):

I got it, its 3.11 :) thanks for the help!